OpenStudy (anonymous):

Hello, can anybody help me figure out what I'm doing wrong on a multiple choose question about slopes. I cannot figure it out for the life of me! I think I may be doing it wrong.. :) Choose the correct slope of the line that passes through the points (−6, −6) and (−3, 1). -7/3 3/7 7/3 -3/7

OpenStudy (anonymous):

from 1st co-ord to 2nd x increases by 3 y increases by 7 gradient = y/x = 7/3

OpenStudy (anonymous):

the slope is 7/3 take y2-y1/x2-x1 you can check by plotting to see if the sign is right. (1- -6)/(-3- -6) (1+6)/(-3+6) 7/3

OpenStudy (anonymous):

Oh, I see now! Thanks!